A pizza shop claims its average home delivery time is 33 minutes. A sample of 35 deliveries had a sample average of 35.9 minuHow would you know to choose a. in part a? As well for part b, how??
Answer #1

Solutions Grent 33 41333 6 = 6.7 n=35 X=35.g PCX235-3) = F-PX 359) =1-P(X-4 , 35:9-33 Sun 6.7/35 I-P(Z <2.5606) = -0.9947 P X6 Consbuct the symmetrical internal that includes 95% of the sample mean if the towe population mean is 33 minutes c. I = (Mt